OpenStudy (kingj):

What is the sum of a 23–term arithmetic sequence where the first term is 9 and the last term is 185? 1,472 1,725 2,231 2,478

OpenStudy (anonymous):

Good question. Let's think for a minute. If we have an arithmetic sequence, we add the same thing over and over. So\[185=9+22k\]if k is that number we add over and over.

OpenStudy (anonymous):

Looks like we add eight repeatedly.

OpenStudy (kingj):

soo what would be the answer?

OpenStudy (anonymous):

Looks like you are only interested in the answer, than how to actually do the question.

OpenStudy (kingj):

I need the answer so I know i did it right

OpenStudy (anonymous):

So the total is\[9+\sum_{1}^{22}9k=9(1+\sum_{1}^{22}k)=9(1+22*23/2)=9(254)=2286\]

OpenStudy (anonymous):

Kay, sry for being offensive, just that this site has too many pple just looking for answers only.

OpenStudy (kingj):

its ok lol

OpenStudy (anonymous):

Doesn't look like my answer meets the specifications of the multiple choice problem. Anybody see where I messed up?

OpenStudy (anonymous):

I found the mistake. Just a second.

OpenStudy (anonymous):

Sum of an arithmetic progression= n/2 ( first term + Nth term) =23/2 (9+185)= 2231

OpenStudy (anonymous):

the mistake lies in the 22 * (23/2). that makes no sense.

OpenStudy (anonymous):

\[9+\sum_{1}^{22}(9+8k)=9+(22)9+8(22*23/2)=\]

OpenStudy (anonymous):

22*23/2 is the sum of the first 22 numbers. That's a formula.

OpenStudy (anonymous):

equals 23(9) + 253(8)=207+2024=2231

OpenStudy (anonymous):

Darn, it's been a long time since I did one of these.....

OpenStudy (anonymous):

My original mistake was to leave out the nine in the terms that were being summed. In addition, I calculated adding nine each time out of not paying attention to what I was doing. Hope I was helpful.

OpenStudy (anonymous):

tiaph, your solution is much more elegant than mine. I hadn't done one like this in a long time and tried brute force. Your solution is elegant.